An economy’s capital stock was $46.3 billion at the end of 2016, $46.6 billion at the end of 2017, and $47.0 billion at the end of 2018. Depreciation in 2017 was $2.4 billion, and gross investment during 2018 was $2.8 billion (all in 2007 dollars).

  1. Calculate net investment and gross investment during 2017.
  2. Calculate depreciation and net investment during 2018.
1
Expert's answer
2022-02-23T12:23:44-0500

"Net Investment_{2017}=C.S_{2017}-C.S_{2016}"

"=4.66-4.63=0.03 billions"

Where C.S is capital stock.

"Gross Investment_{2017}=Net Investment+ Depreciation"

"=0.03+2.4=2.43 billions"

"Net Investment_{2018}=4.7-4.66=0.04 billions"

"Depreciation= Gross Investment- Net Investment"

"=3.8-0.04= 2.76 Billions"